OpenStudy (anonymous):

If you have 3 6-sided dice. The first die has numbers 1-5; with 2 twice. The second die has numbers 3-6, with numbers 4 three times, The third die has only odd numbers from 1-6, with one twice and three thrice. Die one is weighted to where the number 1 is 2% more likely to appear than the number 3 on die 2. Die two is weighted to where 5 is 5% more likely to appear than 5 on die three. Die three is weighted where 3 is 4% less likely to appear than the prob of 2 on die 1 and 4 on die 2. What are the chances of 3 appearing on all 3 die?

OpenStudy (anonymous):

Die I: 1, 2, 2, 3, 4, 5 Die II: 3, 4, 4, 4, 5, 6 Die III: 1, 1, 3, 3, 3, 5 -------------------------------------------------- Die I Prop alpha (unchanged): 1: (1/6) 2: (2/6) = (1/3) 3: (1/6) 4: (1/6) 5: (1/6) 6: (0/6) Die II Prop alpha (unchanged): 1: (0/6) 2: (0/6) 3: (1/6) 4: (3/6) = (1/2) 5: (1/6) 6: (1/6) Die III Prop alpha (unchanged): 1: (2/6) = (1/3) 2: (0/6) 3: (3/6) = (1/2) 4: (0/6) 5: (1/6) 6: (0/6) -------------------------------------------------- Die I Prop. beta (altered, final) 1: 0.17 2: (2/6) = (1/3) 3: (1/6) 4: (1/6) 5: (1/6) 6: (0/6) Die II Prop. beta (altered, final) 1: (0/6) 2: (0/6) 3: (1/6) 4: (3/6) = (1/2) 5: 0.175 6: (1/6) Die III Prop. beta (altered, final) 1: (2/6) = (1/3) 2: (0/6) 3: 0.16 4: (0/6) 5: (1/6) 6: (0/6) probability(2 on die 1 and 4 on die 2) = (1/3)(1/2) = (1/6) -------------------------------------------------------- probability(3 on die 1, 3 on die 2, and 3 on die 3) = (1/6)(1/6)(0.16) = 0.004-bar --------------------------------------------------------------- probability(odd on die 1, even on die 2, odd on die 3) = probability(1 or 3 or 5 on die 1)probability(2 or 4 or 6 on die 2) *probability(1 or 3 or 5 on die 3)

OpenStudy (anonymous):

let probability(1 or 3 of 5 on die 1) = p(a) let probability(2 or 4 or 6 on die 2) = p(b) let probability(1 or 3 or 5 on die 3) = p(c) p(a) = 0.17+(1/6)+(1/6) = 0.503-bar p(b) = (0/6)+(1/2)+(1/6) = (2/3) p(c) = (1/3) +0.16+(1/6) = 0.66 p(a) and p(b) and p(c) = (0.503-bar)(2/3)(0.66) = 0.22146

OpenStudy (anonymous):

@satellite73, could you just see if the work seems right?